# A Survey on 3D Gaussian Splatting

### v7

@@ 5.4 Avatar
Avatars, the digital representations of users in virtual spaces, bridge physical and digital realms, enabling immersive interaction, identity expression, and remote collaboration. Spanning entertainment (gaming, virtual influencers), enterprise (AI agents, virtual meetings), healthcare, and education, they underpin metaverse economies. Advances in AR and VR amplify their role in redefining social, industrial, and creative landscapes.
@@
3D GS has emerged as a powerful tool for human avatar reconstruction, primarily advancing along two directions: full-body modeling and head-centric modeling. For full-body avatars [241, 242, 144, 146, 243, 244, 245, 246, 145, 139, 247, 248, 249, 147, 250, 251, 252], the current methods typically anchor 3D Gaussians in a canonical space and deform them via parametric body models (e.g., SMPL) or cage-based rigging to model dynamic motions. These approaches adopt a hybrid deformation strategy: linear blend skinning handles rigid skeletal transformations such as joint rotations, while pose-conditioned deformation fields account for secondary non-rigid effects like muscle jiggles. For head avatars [149, 150, 253, 23, 254, 255, 148, 151, 256], the emphasis shifts to modeling intricate facial expressions, fine-grained geometry (e.g., wrinkles, hair [257]), and dynamic speech-driven animations. Techniques mainly combine parametric morphable face models (e.g., FLAME) with deformable 3D Gaussians, employing diffusion strategies and expression-aware deformation fields to disentangle rigid head poses from non-rigid facial movements. Both directions exploit the speed advantage and editability of 3D GS to enable efficient training, real-time rendering, and precise control over deformations, while addressing challenges in cross-frame correspondence, topology flexibility, and multi-view consistency.
@@
Reconstruction in challenging scenes (e.g., occlusions, sparse single-view inputs, or loose clothing) and enhancing avatar interactivity represent critical challenges and opportunities. Parametric model-free methods, which bypass predefined priors by learning skinning weights directly from data, show promise for such scenarios. Complementary to this, generative models can mitigate ambiguities inherent in underconstrained settings. Further integrating physics-based constraints might bridge the gap between static reconstructions and responsive, lifelike interactions, unlocking applications in AR, embodied AI, etc.
5.5 Endoscopic Scene Reconstruction
Surgical 3D reconstruction represents a fundamental task in robot-assisted minimally invasive surgery, aimed at enhancing intraoperative navigation, preoperative planning, and educational simulations through precise modeling of dynamic surgical scenes. Pioneering the integration of dynamic radiance fields into this domain, recent advancements have focused on surmounting the inherent challenges of single-viewpoint video reconstructions such as occlusions by surgical instruments and sparse viewpoint diversity within the confined spaces of endoscopic exploration [258, 259, 260]. Despite the progress, the call for high fidelity in tissue deformability and topological variation remains, coupled with the pressing demand for faster rendering to bridge the utility in applications sensitive to latency [152, 153, 154]. This synthesis of immediacy and precision in reconstructing deformable tissues from endoscopic videos is essential in propelling robotic surgery towards reduced patient trauma and AR/VR applications, ultimately fostering a more intuitive surgical environment and nurturing the future of surgical automation and robotic proficiency.
@@ 5.5 Endoscopic Scene Reconstruction
Endoscopic scene reconstruction introduces distinct challenges compared to general dynamic scenes, including sparse training data from limited camera mobility in narrow cavities, frequent tool occlusions obscuring critical regions, and single-view geometry ambiguities. Existing approaches mainly used additional depth guidance to infer the geometry of tissues [152, 153, 154]. For instance, EndoGS [154] integrated depth-guided supervision with spatial-temporal weight masks and surface-aligned regularization terms to enhance the quality and speed of 3D tissue rendering while addressing tool occlusion. EndoGaussian [153] introduced two new strategies: holistic Gaussian initialization for dense initialization and spatiotemporal Gaussian tracking for modeling surface dynamics. Zhao et al[155] argued that these methods suffer from under-reconstruction and proposed to alleviate this problem from frequency perspectives. In addition, EndoGSLAM [156] and Gaussian Pancake [157] devised SLAM systems for endoscopic scenes and showed significant speed advantages.
@@
Advancing endoscopic 3D reconstruction requires targeted efforts in both data and dynamics modeling. Data limitations arise from single-viewpoint videos, which produce ill-posed reconstruction problems due to instrument occlusions and constrained camera mobility, leaving critical tissue regions unobserved. While depth estimators provide temporary workarounds, integrating multi-view camera systems addresses the root cause. In addition, existing datasets often feature truncated sequences (e.g., $4\!\sim\!8s$ in EndoNeRF [258]), which fail to capture prolonged tissue deformation dynamics or complex surgical workflows. Extending temporal coverage to include longer, clinically representative sequences would benefit downstream applications as aforementioned. Modeling limitations persist in current methods, which often represent tissue dynamics at the Gaussian level rather than object- or 3D region-level. This reduces their capacity to encode semantically meaningful anatomical interactions and deserves further explorations.
5.6 Large-scale Scene Reconstruction
@@
Large-scale scene reconstruction is a critical component in fields such as autonomous driving, aerial surveying, and AR/VR, demanding both photorealistic visual quality and real-time rendering capabilities. Before the emergence of 3D GS, the task has been approached using NeRF based methods, which, while effective for smaller scenes, often fall short in detail and rendering speed when scaled to larger areas (e.g., over 1.5 $km^{2}$). Though 3D GS has demonstrated considerable advantages over NeRFs, the direct application of 3D GS to large-scale environments introduces significant challenges. 3D GS requires an immense number of Gaussians to maintain visual quality over extensive areas, leading to prohibitive GPU memory demands and considerable computational burdens during rendering. For instance, a scene spanning 2.7 $km^{2}$ may require over 20 million Gaussians, pushing the limits of even the most advanced hardware (e.g., NVIDIA A100 with 40GB memory) [163].
@@
To address the highlighted challenges, researchers have made significant strides in two key areas: i) For training, a divide-and-conquer strategy [163, 164, 165, 162] has been adopted, which segments a large scene into multiple, independent cells. This facilitates parallel optimization for expansive environments. With the same spirit, Zhao et al[161] proposed a distributed implementation of 3D GS training. An additional challenge lies in maintaining visual quality, as large-scale scenes often feature texture-less surfaces that can hamper the effectiveness of optimization such as Gaussian initialization and density control (Sec. 3.2). Enhancing the optimization algorithm presents a viable solution to mitigate this issue [44, 164]. ii) Regarding rendering, the adoption of the Level of Details (LoD) technique from computer graphics has proven instrumental. LoD adjusts the complexity of 3D scenes to balance visual quality with computational efficiency. Current implementations involve feeding only the essential Gaussians to the rasterizer [164], or designing explicit LoD structures like the Octree [165] and hierarchy [162]. Furthermore, integrating extra input modalities like LiDAR can further enhanced the reconstruction process [158, 159, 160].
@@
One prominent challenge in large-scale scene reconstruction lies in handling sparse or incomplete capture data, which can be mitigated through few-shot adaptation schemes (see Sec. 4.1) or generalizable priors (see “learning physical priors from large-scale data” in Sec. 7). Meanwhile, memory and computational bottlenecks can be addressed via distributed learning strategies [161], such as parameter partitioning across GPU clusters and parallel batched multi-view optimization.
@@
TABLE I: Comparison of localization methods (§6.1) on Replica [261] (static scenes), in terms of absolute trajectory error (ATE, cm). (The three best scores are marked in red, blue, and green, respectively. These notes also apply to the other tables.)
@@
iMAP [262] [ICCV21] 3.12 2.54 2.31 1.69 1.03 3.99 4.05 1.93 2.58
Vox-Fusion [263] [ISMAR22] 1.37 4.70 1.47 8.48 2.04 2.58 1.11 2.94 3.09
NICE-SLAM [264] [CVPR22] 0.97 1.31 1.07 0.88 1.00 1.06 1.10 1.13 1.06
ESLAM [265] [CVPR23] 0.71 0.70 0.52 0.57 0.55 0.58 0.72 0.63 0.63
Point-SLAM [266] [ICCV23] 0.61 0.41 0.37 0.38 0.48 0.54 0.69 0.72 0.52
Co-SLAM [267] [CVPR23] 0.70 0.95 1.35 0.59 0.55 2.03 1.56 0.72 1.00
Gaussian-SLAM [114] [arXiv] 3.35 8.74 3.13 1.11 0.81 0.78 1.08 7.21 3.27
GSSLAM [113] [CVPR24] 0.47 0.43 0.31 0.70 0.57 0.31 0.31 3.20 0.79
GS-SLAM [111] [CVPR24] 0.48 0.53 0.33 0.52 0.41 0.59 0.46 0.70 0.50
SplaTAM [112] [CVPR24] 0.31 0.40 0.29 0.47 0.27 0.29 0.32 0.55 0.36
5.7 Physics
The simulation of complex real-world dynamics, such as seed dispersal or fluid motion, is pivotal for applications spanning virtual reality, animation, and scientific modeling, where realism hinges on accurate physical behavior. Advances in video diffusion models have driven progress in 4D content generation, yet these methods might produce visually plausible results that violate fundamental physical laws. 3D GS emerges as a promising solution by embedding physical constraints and properties into scene representations, enabling both visually convincing and physically coherent simulations.
Existing methods differ in how they formulate and integrate physics-based priors into their frameworks. The most common approach is employing physics simulation engines (e.g., MLS-MPM [268]) to guide the dynamics generation. The material point method [268] and position based dynamics [269] — numerical methods used in computer graphics for simulating deformations in materials like fluids, granular media, and fracturing solids — have been extensively explored by the community through various customizations [166, 167, 168, 143, 169, 170, 21, 171]. Analytical material models, such as mass-spring systems, have also demonstrated success in approximating deformations by explicitly encoding material properties into 3D Gaussians [172]. Across these methods, 3D Gaussians are treated as discrete particles (with one exception [173] using a continuous representation) and serve as computational units within the chosen simulator. Unknown material properties or physical parameters are typically learned through video-based supervision from conditional generative models.
6 Performance Comparison
TABLE II: Collection of representative datasets for 3D GS. Here PC represents point clouds.
     Name Type # Sample Task
Tanks&Temples [270] [TOG17] RGB 14 Novel View Synthesis
RealEstate10K [271] [TOG18] RGB 1,000
DeepBlending [272] [TOG18] RGB 19
LLFF [273] [TOG19] RGB 8
NeRF [12] [ECCV20] RGB 8
ACID [274] [ICCV21] RGB 700+
Mip-NeRF 360 [8] [CVPR22] RGB 9
TUM RGB-D [275] [IROS12] RGB-D 39 Robotics
KITTI [276] [CVPR12] RGB-D&PC 11
ScanNet [277] [CVPR17] RGB-D 1,513
Replica [261] [arXiv19] RGB-D 18
Waymo [278] [CVPR20] RGB-D&PC 1,150
nuScenes [279] [CVPR20] RGB-D&PC 1,000
RLBench [280] [RA-L20] RGB 100
Robomimic [281] [CoRL22] RGB 800
D-NeRF [184] [CVPR21] RGB 8 Dynamic Scene Reconstruction
HyperNeRF [185] [TOG21] RGB 6
NeRF-DS [282] [CVPR23] RGB 8
CoNeRF [283] [CVPR22] RGB 7 Generation and Editing
SPIn-NeRF [284] [CVPR23] RGB 10
Tensor4D [285] [CVPR23] RGB 4
OmniObject3D [286] [CVPR23] 3D Object 6,000
Objaverse [287] [CVPR23] 3D Object 800K+
People-Snapshot [288] [CVPR18] RGB 24 Avatar
VOCASET [289] [CVPR19] RGB 12
THUman [290] [ICCV19] RGB 200
THUman2.0 [291] [CVPR21] RGB-D 500
ZJU-Mocap [292] [CVPR21] RGB 9
H3DS [293] [ICCV21] RGB 23
THUman3.0 [294] [TPAMI22] 3D Scan 20
SCARED [295] [MICCAI19] RGB-D 9 Medical
EndoNeRF [258] [MICCAI22] RGB 2
X3D [296] [CVPR24] X-ray 15
CityNeRF [297] [ECCV22] RGB 12 Large-scale Reconstrction
Waymo Block-NeRF [298] [CVPR22] RGB&PC 1
UrbanBIS [299] [SIGGR23] RGB&PC 6
GauU-Scene [160] [arXiv24] RGB&PC 1

### v8

@@ 5.4. Avatar
Avatars, the digital representations of users in virtual spaces, bridge physical and digital realms, enabling immersive interaction, identity expression, etc. Spanning entertainment (gaming, virtual influencers), enterprise (AI agents, virtual meetings), healthcare, and education, they underpin metaverse economies. Advances in AR and VR amplify their role in redefining social, industrial, and creative landscapes.
@@
3D GS has emerged as a powerful tool for human avatar reconstruction, primarily advancing along two directions: full-body modeling and head-centric modeling. For full-body avatars (Jena et al., 2023; Ye et al., 2023; Li et al., 2024k; Lei et al., 2024a; Moreau et al., 2024; Kocabas et al., 2024; Abdal et al., 2024; Zheng et al., 2024b; Hu and Liu, 2024; Jiang et al., 2024a; Hu et al., 2024c; Pang et al., 2024; Qian et al., 2024b; Yuan et al., 2024; Jung et al., 2023; Li et al., 2023b, 2024h), the current methods typically anchor 3D Gaussians in a canonical space and deform them via parametric body models (e.g., SMPL) or cage-based rigging to model dynamic motions. These approaches adopt a hybrid deformation strategy: linear blend skinning handles rigid skeletal transformations such as joint rotations, while pose-conditioned deformation fields account for secondary non-rigid effects like muscle jiggles. For head avatars (Qian et al., 2024a; Dhamo et al., 2023; Xiang et al., 2023; Saito et al., 2024; Chen et al., 2023b; Zhao et al., 2024a; Zhou et al., 2024d; Li et al., 2024j; Rivero et al., 2024), the emphasis shifts to modeling intricate facial expressions, fine-grained geometry (e.g., wrinkles, hair (Luo et al., 2024)), and dynamic speech-driven animations. Techniques mainly combine parametric morphable face models (e.g., FLAME) with deformable 3D Gaussians, employing diffusion strategies and expression-aware deformation fields to disentangle rigid head poses from non-rigid facial movements. Both directions exploit the speed advantage and editability of GS to enable fast training, real-time rendering, and precise control over deformations, while addressing challenges in cross-frame correspondence, topology flexibility, and multi-view consistency.
@@
Reconstruction in challenging scenes (e.g., occlusions, sparse single-view inputs, or loose clothing) and enhancing avatar interactivity represent critical challenges and opportunities. Parametric model-free methods, which bypass predefined priors by learning skinning weights directly from data, show promise for such scenarios. Complementary to this, generative models can mitigate ambiguities inherent in underconstrained settings. Further integrating physics-based constraints might bridge the gap between static reconstructions and responsive, lifelike interactions, unlocking applications in metaverse, embodied AI, etc.
5.5. Endoscopic Scene Reconstruction
Endoscopic Scene reconstruction is pivotal in robot-assisted minimally invasive surgery, enhancing intraoperative navigation, planning, and simulation through accurate modeling of dynamic scenes. Recent advances integrate dynamic radiance fields to address challenges like instrument occlusions and sparse viewpoints in endoscopic videos. However, achieving high fidelity in tissue deformation and topological variation, alongside real-time rendering for latency-sensitive applications, remains critical (Huang et al., 2024b; Liu et al., 2024b; Zhu et al., 2024c).
@@ 5.5. Endoscopic Scene Reconstruction
This task introduces distinct challenges compared to general dynamic scenes, including sparse training data from limited camera mobility in narrow cavities, tool occlusions obscuring critical regions, and single-view geometry ambiguities. Existing approaches mainly used additional depth guidance to infer the geometry of tissues (Huang et al., 2024b; Liu et al., 2024b; Zhu et al., 2024c). For instance, EndoGS (Zhu et al., 2024c) integrated depth-guided supervision with spatial-temporal weight masks and surface-aligned regularization terms to enhance the quality and speed of 3D tissue rendering while addressing tool occlusion. EndoGaussian (Liu et al., 2024b) introduced two new strategies: holistic Gaussian initialization for dense initialization and spatiotemporal Gaussian tracking for modeling surface dynamics. Zhao et al. (Zhao et al., 2024d) argued that these methods suffer from under-reconstruction and proposed to alleviate this problem from frequency perspectives. In addition, EndoGSLAM (Wang et al., 2024c) and Gaussian Pancake (Bonilla et al., 2024) devised SLAM systems for endoscopic scenes and showed significant speed advantages.
@@
Advancing endoscopic 3D reconstruction requires targeted efforts in both data and dynamics modeling. Data limitations arise from single-viewpoint videos, which produce ill-posed reconstruction problems due to instrument occlusions and constrained camera mobility, leaving critical tissue regions unobserved. While depth estimators provide temporary workarounds, integrating multi-view camera systems addresses the root cause. In addition, existing datasets often feature truncated sequences (e.g., $4\!\sim\!8s$ in EndoNeRF (Wang et al., 2022)), which fail to capture prolonged tissue deformation dynamics or complex surgical workflows. Extending temporal coverage to include longer, clinically representative sequences would benefit downstream applications as aforementioned. Modeling limitations persist in current methods, which often represent tissue dynamics at the Gaussian level rather than object- or 3D region-level. This reduces their capacity to encode semantically meaningful anatomical interactions and deserves further explorations.
5.6. Large-scale Scene Reconstruction
@@
Large-scale scene reconstruction is a critical component in fields such as autonomous driving, aerial surveying, and AR/VR, demanding both photorealistic visual quality and real-time rendering capabilities. Before the emergence of 3D GS, the task has been approached using NeRF based methods, which, while effective for smaller scenes, often fall short in detail and rendering speed when scaled to larger areas (e.g., over 1.5 $km^{2}$). Though 3D GS has demonstrated considerable advantages over NeRFs, the direct application of 3D GS to large-scale environments introduces significant challenges. 3D GS requires an immense number of Gaussians to maintain visual quality over extensive areas, leading to prohibitive GPU memory demands and considerable computational burdens during rendering. For instance, a scene spanning 2.7 $km^{2}$ may require over 20 million Gaussians, pushing the limits of even the most advanced hardware (e.g., NVIDIA A100 with 40GB memory) (Liu et al., 2024a).
@@
To address the highlighted challenges, researchers have made significant strides in two key areas: i) For training, a divide-and-conquer strategy (Liu et al., 2024a; Lin et al., 2024b; Ren et al., 2024; Kerbl et al., 2024) has been adopted, which segments a large scene into multiple, independent cells. This facilitates parallel optimization for expansive environments. An additional challenge lies in maintaining visual quality, as large-scale scenes often feature texture-less surfaces that can hamper the effectiveness of optimization such as Gaussian initialization and density control (Sec. 3.2). Enhancing the optimization algorithm presents a viable solution to mitigate this issue (Cheng et al., 2024; Lin et al., 2024b). ii) Regarding rendering, the adoption of the Level of Details (LoD) technique from computer graphics has proven instrumental. LoD adjusts the complexity of 3D scenes to balance visual quality with computational efficiency. Current implementations involve feeding only the essential Gaussians to the rasterizer (Lin et al., 2024b), or designing explicit LoD structures like the Octree (Ren et al., 2024) and hierarchy (Kerbl et al., 2024). Furthermore, integrating extra input modalities like LiDAR can further enhanced the reconstruction process (Wu et al., 2024g, b; Xiong et al., 2024).
@@
One prominent challenge lies in handling sparse or incomplete capture data, which can be mitigated through few-shot adaptation schemes (see Sec. 4.1) or generalizable priors (see “learning physical priors from large-scale data” in Sec. 7). Meanwhile, memory and computational bottlenecks can be addressed via distributed learning strategies (Zhao et al., 2024c), such as parameter partitioning across GPU clusters and parallel batched multi-view optimization.
5.7. Physics
The simulation of complex real-world dynamics, such as seed dispersal or fluid motion, is pivotal for applications spanning VR, animation, and scientific modeling, where realism hinges on accurate physical behavior. Advances in diffusion models have driven progress in 4D content generation, yet these methods might produce visually plausible results that violate fundamental physical laws. 3D GS emerges as a promising solution by embedding physical constraints and properties into scene representations, enabling both visually convincing and physically coherent simulations.
Existing methods differ in how they formulate and integrate physics-based priors into their frameworks. The most common approach is employing physics simulation engines (e.g., MLS-MPM (Hu et al., 2018)) to guide the dynamics generation. The material point method (Hu et al., 2018) and position based dynamics (Müller et al., 2007) — numerical methods used in computer graphics for simulating deformations in materials like fluids, granular media, and fracturing solids — have been extensively explored by the community through various customizations (Xie et al., 2024; Liu et al., 2024e; Borycki et al., 2024; Qiu et al., 2024; Huang et al., 2025; Zhang et al., 2024h; Jiang et al., 2024c; Feng et al., 2024a). Analytical material models, such as mass-spring systems, have also demonstrated success in approximating deformations by explicitly encoding material properties into 3D Gaussians (Zhong et al., 2024). Across these methods, 3D Gaussians are treated as discrete particles (with one exception (Shao et al., 2024a) using a continuous representation) and serve as computational units within the chosen simulator. Unknown material properties or physical parameters are typically learned through video-based supervision from conditional generative models.
Table 1. Collection of representative datasets for 3D GS. Here PC represents point clouds.
       Name Type # Sample Area and Task
Tanks&Temples (Knapitsch et al., 2017) [TOG17] RGB 14 Novel View Synthesis
RealEstate10K (Zhou et al., 2018) [TOG18] RGB 1,000
DeepBlending (Hedman et al., 2018) [TOG18] RGB 19
LLFF (Mildenhall et al., 2019) [TOG19] RGB 8
NeRF (Mildenhall et al., 2020) [ECCV20] RGB 8
ACID (Liu et al., 2021) [ICCV21] RGB 700+
Mip-NeRF 360 (Barron et al., 2022) [CVPR22] RGB 9
TUM RGB-D (Sturm et al., 2012) [IROS12] RGB-D 39 Robotics
KITTI (Geiger et al., 2012) [CVPR12] RGB-D&PC 11
ScanNet (Dai et al., 2017) [CVPR17] RGB-D 1,513
Replica (Straub et al., 2019) [arXiv19] RGB-D 18
Waymo (Sun et al., 2020) [CVPR20] RGB-D&PC 1,150
nuScenes (Caesar et al., 2020) [CVPR20] RGB-D&PC 1,000
RLBench (James et al., 2020) [RA-L20] RGB 100
Robomimic (Mandlekar et al., 2022) [CoRL22] RGB 800
D-NeRF (Pumarola et al., 2021) [CVPR21] RGB 8 Dynamic Scene Reconstruction
HyperNeRF (Park et al., 2021b) [TOG21] RGB 6
NeRF-DS (Yan et al., 2023) [CVPR23] RGB 8
CoNeRF (Kania et al., 2022) [CVPR22] RGB 7 Generation and Editing
SPIn-NeRF (Mirzaei et al., 2023) [CVPR23] RGB 10
Tensor4D (Shao et al., 2023) [CVPR23] RGB 4
OmniObject3D (Wu et al., 2023) [CVPR23] 3D Object 6,000
Objaverse (Deitke et al., 2023) [CVPR23] 3D Object 800K+
People-Snapshot (Alldieck et al., 2018) [CVPR18] RGB 24 Avatar
VOCASET (Cudeiro et al., 2019) [CVPR19] RGB 12
THUman (Zheng et al., 2019) [ICCV19] RGB 200
THUman2.0 (Yu et al., 2021b) [CVPR21] RGB-D 500
ZJU-Mocap (Peng et al., 2021b) [CVPR21] RGB 9
H3DS (Ramon et al., 2021) [ICCV21] RGB 23
THUman3.0 (Su et al., 2022) [TPAMI22] 3D Scan 20
SCARED (Allan et al., 2021) [MICCAI19] RGB-D 9 Medical Systems
EndoNeRF (Wang et al., 2022) [MICCAI22] RGB 2
X3D (Cai et al., 2024b) [CVPR24] X-ray 15
CityNeRF (Xiangli et al., 2022) [ECCV22] RGB 12 Large-scale Reconstrction
Waymo Block-NeRF (Tancik et al., 2022) [CVPR22] RGB&PC 1
UrbanBIS (Yang et al., 2023c) [SIGGR23] RGB&PC 6
GauU-Scene (Xiong et al., 2024) [arXiv24] RGB&PC 1
6. Performance Comparison
In this section, we provide more empirical evidence by presenting the performance of several 3D GS algorithms that we previously discussed. The diverse applications of 3D GS across numerous tasks, coupled with the custom-tailored algorithmic designs for each task, render a uniform comparison of all 3D GS algorithms across a single task or dataset impracticable. For comprehensiveness, we provide a collection of representative datasets in Table 5.7 Following the taxonomy outlined Fig. 2. Due to the limited space, we have chosen several representative tasks for an in-depth performance evaluation. The performance scores are primarily sourced from the original papers, except where indicated otherwise. We also maintain a Github repository for this section.
@@ 6. Performance Comparison
Table 2. Comparison of localization methods on Replica (Straub et al., 2019) (static scenes), in terms of absolute trajectory error (ATE, cm). (The three best scores are marked in red, blue, and green, respectively. These notes also apply to the other tables.)
@@
iMAP (Sucar et al., 2021) [ICCV21] 3.12 2.54 2.31 1.69 1.03 3.99 4.05 1.93 2.58
Vox-Fusion (Yang et al., 2022) [ISMAR22] 1.37 4.70 1.47 8.48 2.04 2.58 1.11 2.94 3.09
NICE-SLAM (Zhu et al., 2022) [CVPR22] 0.97 1.31 1.07 0.88 1.00 1.06 1.10 1.13 1.06
ESLAM (Johari et al., 2023) [CVPR23] 0.71 0.70 0.52 0.57 0.55 0.58 0.72 0.63 0.63
Point-SLAM (Sandström et al., 2023) [ICCV23] 0.61 0.41 0.37 0.38 0.48 0.54 0.69 0.72 0.52
Co-SLAM (Wang et al., 2023b) [CVPR23] 0.70 0.95 1.35 0.59 0.55 2.03 1.56 0.72 1.00
Gaussian-SLAM (Yugay et al., 2023) [arXiv] 3.35 8.74 3.13 1.11 0.81 0.78 1.08 7.21 3.27
GSSLAM (Matsuki et al., 2024) [CVPR24] 0.47 0.43 0.31 0.70 0.57 0.31 0.31 3.20 0.79
GS-SLAM (Yan et al., 2024c) [CVPR24] 0.48 0.53 0.33 0.52 0.41 0.59 0.46 0.70 0.50
SplaTAM (Keetha et al., 2024) [CVPR24] 0.31 0.40 0.29 0.47 0.27 0.29 0.32 0.55 0.36
6.1. Performance Benchmarking: Localization
$\bullet$ Dataset: Replica (Straub et al., 2019) dataset is a collection of 18 highly detailed 3D indoor scenes. These scenes are not only visually realistic but also offer data including dense meshes, high-quality textures, and detailed semantic information for each element. Following (Sucar et al., 2021), three sequences about rooms and five sequences about offices are used for the evaluation.
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve four recent 3D GS based algorithms (Yugay et al., 2023; Matsuki et al., 2024; Keetha et al., 2024; Yan et al., 2024c) and six typical SLAM methods (Zhu et al., 2022; Sandström et al., 2023).
$\bullet$ Evaluation Metric: The root mean square error (RMSE) of the absolute trajectory error (ATE) is a commonly used metric in evaluating SLAM systems (Sturm et al., 2012), which measures the root mean square of the Euclidean distances between the estimated and true positions over the entire trajectory.
$\bullet$ Result: As shown in Table 6, the recent 3D Gaussians based localization algorithms have a clear advantage over existing NeRF based dense visual SLAM. For example, SplaTAM (Keetha et al., 2024) achieves a trajectory error improvement of $\sim$50%, decreasing it from 0.52cm to 0.36cm compared to the previous state-of-the-art (SOTA) (Sandström et al., 2023). We attribute this to the dense and accurate 3D Gaussians reconstructed for scenes, which can handle the noise of real sensors. This reveals that effective scene representations can improve the accuracy of localization tasks.
6.2. Performance Benchmarking: Static Scenes